OpenStudy (anonymous):

Hexagon A is a regular hexagon. The total length of all the sides of the hexagon is 36 inches. Hexagon A is dilated about its center to create Hexagon B. The length of each side of Hexagon B is 1 & 1/2 inches. By what factor was Hexagon A dilated to create Hexagon B?

jimthompson5910 (jim_thompson5910):

1 & 1/2 = 1 + 1/2 = 1 + 0.5 = 1.5 Now divide 1.5 by 36 to get your answer

OpenStudy (anonymous):

the answer is 4 right? :)

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

oh wait, not sure why i'm thinking 36 lol

OpenStudy (anonymous):

i mean 1/4

jimthompson5910 (jim_thompson5910):

each side of A is 6, not 36, my bad 1.5/6 = 0.25 = 1/4

jimthompson5910 (jim_thompson5910):

so yes

OpenStudy (anonymous):

thanks :)

jimthompson5910 (jim_thompson5910):

yw